How can we represent the number of radial nodes and angular nodes for d orbital

Grade:12

1 Answers

meghana
6 years ago
there is a formula for radial and angular nodes.no.of radial nodes=n-l-1 and no.of angular nodes=l.For d orbital,l=2 for example 3d orbital no.of radial nodes=3-2-1=0 and no.of angular nodes=2